in library shelf there are 5 novels,3 poems and dictionary.if three books are picked at random then how many ways this can be done if
a)there is no restriction.
b)A dicttionary is selected.
c)2 novels and 1 poem are selected.

Answer by rothauserc(4718)

5 novels,3 poems and dictionary = 9 books
:
a) 9C3 = 9! / (3! * (9-3)!) = (9*8*7) / (3*2) = 3*4*7 = 84 ways
:
b) 1 * 8C2 = 28 ways
:
c) 5C2 * 3C1 = 10 * 3 = 30 ways
: